If 8 people can pack 60 cartons in 2 hours how many carton will 3 people pack in 2 hours
Answer by stanbon(57442)

If 8 people can pack 60 cartons in 2 hours how many carton will 3 people pack in 2 hours
---------------------
# of people and time are inversely related; people and cartons are directly related
p = kc/t
---
Find "k":
8 = k*60/2
k= 8/30 = 4/15
---
So p = (4/15)c/t
----
how many carton will 3 people pack in 2 hours
3 = (4/15)c/2
(4/15)c = 6
c = 15*6/4
c = 22.5 cartons
==============================
Cheers,
Stan H
Reply to stanbon@comcast.net